3. Variable Costing Income Statement

On April 30, the end of the first month of operations, Joplin Company prepared the following income statement, based on the absorption costing concept:

Joplin Company
Absorption Costing Income Statement
For the Month Ended April 30
Sales (6,400 units) $198,400
Cost of goods sold:
Cost of goods manufactured (7,500 units) $165,000
Inventory, April 30 (1,100 units) (24,200)
Total cost of goods sold (140,800)
Gross profit $57,600
Selling and administrative expenses (33,520)
Operating income $24,080

If the fixed manufacturing costs were $37,950 and the fixed selling and administrative expenses were $16,420, prepare an income statement according to the variable costing concept. Round all final answers to whole dollars.

Joplin Company
Variable Costing Income Statement
For the Month Ended April 30
$
Variable cost of goods sold:
$
$
$
Fixed costs:
$
$

4. Change in Sales Mix and Contribution Margin

Head Pops Inc. manufactures two models of solar-powered, noise-canceling headphones: Sun Sound and Ear Bling models. The company is operating at less than full capacity. Market research indicates that 22,800 additional Sun Sound and 25,100 additional Ear Bling headphones could be sold. The operating income by unit of product is as follows:

Sun Sound
Headphones
Ear Bling
Headphones
Sales price $32.20 $50.20
Variable cost of goods sold (18.00) (28.10)
Manufacturing margin $14.20 $22.10
Variable selling and administrative expenses (6.40) (10.00)
Contribution margin $7.80 $12.10
Fixed manufacturing costs (2.90) (4.50)
Operating income $4.90 $7.60

Prepare an analysis indicating the increase or decrease in total profitability if 22,800 additional Sun Sound and 25,100 additional Ear Bling headphones are produced and sold, assuming that there is sufficient capacity for the additional production. Round your per unit answers to two decimal place.

Head Pops Inc.
Analysis
Sun Sound Headphones Ear Bling Headphones
Unit volume increase
x Contribution margin per unit $ $
Increase in profitability $ $
Answer #1
Sales $      1,98,400
Variable Expenses
Variable Cost of Goods Manufactured $      1,27,050 =165000-37950
Less Ending Inventory $         18,634 =127050/7500*1100
Variable Cost of Goods Sold $      1,08,416
Variable Selling and administrative expenses $         17,100 =33520-16420
Total Variable expenses $      1,25,516
Contribution Margin $          72,884
Fixed Costs:
Cost of Goods Sold $         37,950
Selling and administrative expenses $         16,420
Total Fixed Costs $          54,370
Operating Income $          18,514
Sun Sound Headphones Ear Bling Headphones
Unit volume increase 22800 25100
x Contribution margin per unit $             7.80 $            12.10
Increase in profitability $      1,77,840 $      3,03,710